OpenStudy (smurfy14):

(x-2)over(3) + (x+5)over(2)=1/3 Help!!

OpenStudy (anonymous):

2(x-2)/6 + 3(x+5)/6 = 2/6 (2x-4)/6 + (3x+15)/6 = 2/6 5x+11/6 = 2/6 5x = -9/6 x = -45/6

OpenStudy (anonymous):

2(x-2)/6 + 3(x+5)/6 = 2/6 (2x-4)/6 + (3x+15)/6 = 2/6 5x+11/6 = 2/6 5x = -9/6 x = -45/6

OpenStudy (smurfy14):

thanks!

OpenStudy (anonymous):

ANYTIME

OpenStudy (smurfy14):

but actually the answer is x=-9/5